OpenStudy (anonymous):

tan2(x) = 32sec(x) solve for x between [0,2pi]

OpenStudy (anonymous):

use this identity: \[\tan^2x + 1 = \sec^2x\]

OpenStudy (cwrw238):

do you mean tan^2 x ? the square of tan x?

OpenStudy (cwrw238):

yes eigens got tthe way to go

OpenStudy (anonymous):

\[\sec^2x - 1 = 32secx\]\[\sec^2x -32secx -1 = 0\]let secx = y\[y^2 - 32y -1 =0\] solve the quadratic and you will have possible values of sec(x)

OpenStudy (anonymous):

\[\tan(2x)=\sin(2x)/\cos(2x)= 2\sin(x)\cos(x)/1-2\sin^2(x)\]

OpenStudy (anonymous):

is way simpler to solve the given equation, also rewrite the sec(x) = 1/cos(x)

OpenStudy (anonymous):

i think it is meant to be tan^2(x) , which would suggest the easiest way is to use the identity. otherwise you get sin^2(x) = 32cos(x) and you have to use the identity anyway :)

OpenStudy (anonymous):

\[\tan^2x or \tan^(2x)\]

OpenStudy (anonymous):

i think it is probably meant to be tan^2(x)

OpenStudy (anonymous):

tan^2, sorry.

OpenStudy (anonymous):

\[\sec^2x - 32secx - 1 = 0\] is a quadratic \[secx = \frac{32 \pm \sqrt{(1024 +4)}}{2}\]